OpenStudy (anonymous):

The length of a rectangle is three times its width. If the perimeter is at most 112 centimeters, what is the greatest possible value for the width? Write an inequality to model the problem? can anyone help me please?

OpenStudy (anonymous):

Givens: P = 2W + 2L L = 3W P= 112 Substitute 3W into L for the P equation: 112 = 2W + 2(3W) 112 = 2W + 6W 112 = 8W W = 14cm The width is at most 14 cm, so: \[112\ge8W;14\ge W\]

OpenStudy (anonymous):

thank you so much i appreciate that you hella smart